Solve the following system of equations for z and for y: System of Equations: Value of z Value of y y-8+22 y- 56- 6z Solve thNote: Use the orange line (square symbols) to plot the first equation, and use the blue line (circle symbols) to plot the secThe values of p and q that solve these two equations simultaneously can be seen on the graph asthe inverse of the slopes of the two lines the values of the intercepts of the two lines the slopes of the two lines the coor
Answer #1

Given System -56-6z - 56-6乙=8十22 Use this m any of or i we obhan 2 Given System 44+2b- 2o fom i we nawe. 4a: 20-26 (ii) we gep=6_2q 2 plot p=1+q Plot: 10 5 9 2 4It can be seen from the graph

The value of p=8/3 and

Value of q=5/3